# S1-Bench: A Simple Benchmark for Evaluating System 1 Thinking Capability of Large Reasoning Models

### v1

@@ 4.2 Evaluation Metrics
We calculate accuracy metrics for both strict and loose formatting requirements, respectively. We use GPT-4o as the evaluator to assess the correctness of the responses [^8] If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed., with the evaluation prompt in Table 11. For greedy sampling, we directly calculate the accuracy rate, with complete results in Appendix D. For top-p sampling, we utilize two metrics:
@@
Notably, S-Corr represents the upper bound for pass@1 and acc@5 under strict formatting requirements, with L-Corr following the same principle.
4.3 Main Results
Notably, state-of-the-art LRMs, such as DeepSeek-R1 and QwQ-32B, do not demonstrate a distinct advantage in efficiency. In contrast, Sky-T1-32B, which undergoes specific optimizations to mitigate overthinking using SimPO, achieves the highest efficiency. The L-R1-DS 7B/14B/32B models generate longer responses than the DS-R1-7B/14B/32B models, and the former are further post-trained on the latter. This suggests that further post-training may improve reasoning ability for complex tasks, but at the cost of response efficiency for simple tasks. Finally, the s1.1 models generate considerably longer responses than the DeepSeek-R1-Distilled models. Despite both models being trained solely with SFT to acquire long-COT reasoning ability, the DeepSeek-R1-Distilled models use 800K training samples, while the s1.1 models are trained on only 1K. This discrepancy suggests that the smaller training set may lead to superficial imitation of long reasoning patterns, resulting in verbose thinking on simple questions.
Several LRMs exhibit under-accuracy on simple questions.
Our observations reveal that, despite employing deep reasoning, most LRMs tend to exhibit lower accuracy on simple questions compared to traditional LLMs. For example, DS-R1-1.5B and EXAONE-2.4B achieve just above 50% acc@k. Moreover, accuracy decreases with smaller model size. Finally, many LRMs struggle with robust correctness in top-p sampling, where acc@k is significantly lower than pass@1. This issue is particularly pronounced in smaller LRMs. For instance, DS-R1-1.5B achieved 81.47% pass@1 but only 54.50% acc@k.
5 Efficiency Analysis
Figure 4: (a) Comparison of Initial and Additional Thinking Costs for Each LRM. (b) Distribution of Solution Rounds for Each LRM.
5.1 ART Analysis of Question Types
To better understand the efficiency differences of LRMs across question types, we analyze $\mathtt{ART}$ across 4 main categories, 28 subcategories, and two languages, yielding the following insights.
LRMs exhibit the most significant increase in ART for instruction-following questions and tend to over-explore when the solution space is vast.
Although small LLMs generate the most concise responses for instruction-following questions across all main categories in S1-Bench, LRMs produce responses that are 109.7× longer in Chinese and 62.7× longer in English, making them the longest response type across all categories. We further analyze the subcategories of instruction-following questions. As shown in Figure 9, $\mathtt{ART}$ is notably longer in the subcategories of length constraints, character constraints, and sentence constraints. These three question types share a similar characteristic: their correctness is verifiable, but the solution space is vast. We find that, although the model quickly identifies a correct answer, it becomes trapped in the search space, continually exploring alternatives and failing to stop in time. This phenomenon is more pronounced in families with higher $\mathtt{ART}$, such as s1.1 and the EXAONE.
LRMs also demonstrate a significant increase in ART for reasoning tasks.
In Table 3, $\mathtt{ART}$ for reasoning questions also increases significantly, and Figure 9 further corroborates this phenomenon in $\mathtt{ART}$ in various reasoning subcategories. One possible explanation for this is that reasoning questions align closely with the post-training data distribution of LRMs, thereby further stimulating the long-chain reasoning behavior of LRMs.
LRMs consistently exhibit inefficiency in both Chinese and English.
In table 3, most LRMs show higher $\mathtt{ART}$ on English questions, while QwQ-32B and Hunyuan-T1 exhibit comparable $\mathtt{ART}$ across Chinese and English.
5.2 Solution Analysis in Thinking Processes
To better understand the causes of inefficiency in LRMs on the S1-Bench, we analyze the thinking process where the final answer is correct and the format is strictly correct and non-empty [^9] We selected this subset because incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed or empty thinking processes cannot be precisely extracted.. We begin by segmenting each thinking process into several solutions, each defined as a point at which LRMs explicitly arrive at a conclusion that directly aligns with the correct answer. The segmentation process is performed by DeepSeek-v3, with prompts detailed in Table 12. We then compute the average initial thinking cost for LRMs. For each sample, if the thinking process contains at least one solution, the cost is defined as the tokens counts in the first solution. If no clear and correct solution is provided, the cost is the total token counts in the thinking process. The reported value is the average across all selected samples. We find that:
The step-by-step reasoning in the initial thinking process is not the primary cause of the inefficiency in LRMs.
As shown in Figure 4 (a), although the total counts of thinking tokens varies considerably across LRMs, their initial thinking costs are similar and account for only a small fraction of the total.
Generating unnecessary solution rounds after reaching the correct answer is one of the reasons for the inefficiency of LRMs.
We further examine the distribution of solution rounds among various LRMs on the S1-Bench (Figure 4 (b)) and find that models with longer thinking processes tend to produce excessive solution rounds, repeatedly re-verifying simple problems that have already been solved. This redundant verification significantly contributes to computational inefficiency.
5.3 Redundancy in Thinking Processes
Information redundancy increases as reasoning sequences increase. We conduct a similarity analysis to analyze how information redundancy in the thinking processes of LRMs changes as reasoning sequences increase. Specifically, we first divide the complete thinking process into k equal-length segments [^10] We set k=15, changing its value does not affect the conclusions.. Then, we encode each segment using the all-MiniLM-L6-v2 model [^11] https://huggingface.co/sentence-transformers/allMiniLM-L6-v2. For each segment, we calculate the cosine similarity with all its preceding segments and use the maximum similarity as a measure of its information redundancy. As shown in Figure 5, information redundancy increases across all four main categories as reasoning sequences increase. Sky-T1-32B shows overall lower similarity, which stems from its shorter thinking process, but still demonstrates an upward trend.
6 Analysis of Errors
Figure 7 explores the relationship between thinking process quality and final answer accuracy. We focus on samples whose format is strictly correct and thinking process is non-empty. For samples with correct final answers, we categorize them based on whether the thinking process contains explicit incorrect conclusions in intermediate steps. For samples with incorrect final answers, we categorize them based on whether the correct answer is mentioned at least once during reasoning. We use DeepSeek-v3 for categorization, with the prompt provided in Table 13. Figure 6 presents the distribution of these four categories in different RLMs. Our analysis yields the following observations: (1) LRMs with lower accuracy often include incorrect intermediate conclusions in their reasoning, even when they ultimately reach correct final answers (light green). (2) Although LRMs sometimes reach the correct answer during reasoning, they may deviate and ultimately produce incorrect conclusions (light red).
7 Simplicity Prejudgement
Figure 7: Counting instances of prejudgment generated by LRMs for simple questions and $\mathtt{ART}$.
Finally, we discover an intriguing phenomenon: LRMs can prejudge certain simple questions. Specifically, we extract the first 50 tokens from each thinking process and use DeepSeek-v3 to identify sentences showing judgment, further extracting judgment words [^12] An example of prejudgment: “This question seems simple”. For prompt references, see Table 14.. Figure 7 displays the frequency of words judging questions as simple, with details provided in the Table 7. Our findings reveal:
LRMs possess the ability to prejudge question simplicity, especially in Chinese.
All LRMs exhibit prejudgment phenomena within their thinking processes, demonstrating an ability to directly assess question difficulty. Additionally, 14 LRMs exhibit significantly stronger prejudgment tendencies in Chinese, with this phenomenon concentrating primarily within the L-R1, DS-R1, QwQ, and Hunyuan model families.
Even with prejudgment, the thinking length of LRMs does not shorten.
As shown in Figure 7, the average $\mathtt{ART}$ of the thinking process exhibiting prejudgment does not decrease. We consider further exploration of this phenomenon as an important direction for future work.
These results suggest LRMs possess an inherent understanding of question difficulty. This opens a novel pathway toward dual-system compatibility for LRMs, which we identify as future work.
8 Conclusion
This paper presents S1-Bench, designed to evaluate LRMs’ performance on simple questions suited for system 1 thinking. Experiments reveal significant overthinking and performance degradation by LRMs on S1-Bench, highlighting the gap between current LRMs and dual-system compatibility.
Appendix A Subcategories in S1-Bench
Figure 8 shows the pie chart distribution of 28 subcategories in S1-Bench. For more details on the subcategories, please refer to Table 15,16.
Appendix B Baseline Model Details
Model Model ID URL
Model ID Abbreviation Base Model Alg. Size
DeepSeek
DeepSeek-R1-Distill-Qwen-7B DS-R1-7B Qwen2.5-Math-7B SFT 800K
DeepSeek-R1-Distill-Llama-8B DS-R1-8B Llama-3.1-8B SFT 800K
DeepSeek-R1-Distill-Qwen-14B DS-R1-14B Qwen2.5-14B SFT 800K
DeepSeek-R1-Distill-Qwen-32B DS-R1-32B Qwen2.5-32B SFT 800K
DeepSeek-R1 DS-R1 DeepSeek-V3-0324 SFT&RL 800K&–
Qwen
QwQ-32B QwQ-32B Qwen2.5-32B
qihoo360
Light-R1-7B-DS L-R1-7B-DS DeepSeek-R1-Distill-Qwen-7B SFT 3K
Light-R1-32B-DS L-R1-32B-DS DeepSeek-R1-Distill-Qwen-32B SFT 3K
Light-R1-32B L-R1-32B Qwen2.5-32B-Instruct SFT&DPO 73K&–
simplescaling
s1.1-7B s1.1-7B Qwen2.5-7B-Instruct SFT 1K
s1.1-14B s1.1-14B Qwen2.5-14B-Instruct SFT 1K
s1.1-32B s1.1-32B Qwen2.5-32B-Instruct SFT 1K
LG AI Research
NVIDIA
NovaSky
Table 5: The open-source LRMs details evaluated for S1-Bench.
Table 4 presents the abbreviations, IDs, and URLs of LLMs used in this paper. Table 5 displays the abbreviations, IDs, URLs, organizations, algorithms, and training data volumes of open-source LRMs evaluated in this study.
Appendix C Format Details
@@
(standard) (special) (number) process answer
Strict 100 $\surd$ 1 $\surd$ $\surd$
@@
Loose (w/o Strict) 200 $\surd$ 1 $\surd$ $\times$
@@
Error 300 $\surd$ $\surd$ $\geq$1 $\times$ $\times$
@@
Table 6: Format types.
Words Count
{CJK}UTF8gkai简单 466
straightforward 231
simple 118
{CJK}UTF8gkai不难 72
{CJK}UTF8gkai基础的 44
basic 24
{CJK}UTF8gkai直接 15
{CJK}UTF8gkai常见 10
{CJK}UTF8gkai基础 10
{CJK}UTF8gkai不复杂 7
simplest 7
pretty straightforward 5
easy 5
{CJK}UTF8gkai基本的 5
{CJK}UTF8gkai基本 3
Table 7: Top 15 phrases with prejudgment phenomenon, with counts summed across 22 LRMs.
Model Loose Format Error Format
Strict Format
100 101 200 201 202 203 204 205 206 207 300 301
QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Sky-T1-32B 95.26 0.00 0.62 0.09 0.19 0.00 0.28 0.00 0.00 3.03 0.00 0.52
DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-70B 99.91 0.00 0.09 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-14B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
DS-R1-7B 99.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.76
DS-R1-1.5B 97.58 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.42
L-R1-32B 95.07 0.00 0.00 0.00 0.00 0.00 0.81 0.00 0.00 3.03 0.00 1.09
L-R1-7B-DS 99.67 0.00 0.05 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
s1.1-32B 99.53 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.43
s1.1-14B 97.39 0.00 0.00 0.14 0.00 0.00 0.24 0.00 0.00 0.00 0.00 2.23
s1.1-7B 88.96 0.00 0.00 7.96 0.09 0.00 0.00 0.00 0.00 0.09 0.00 2.89
Table 8: Format types rates in top-p sampling setting.
Model Loose Format Error Format
Strict Format
100 101 200 201 202 203 204 205 206 207 300 301
QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Sky-T1-32B 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.47 0.00 0.24
DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-70B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-14B 99.76 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
DS-R1-7B 97.87 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.13
DS-R1-1.5B 91.94 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.06
L-R1-32B 85.55 0.24 0.00 0.24 0.71 0.24 0.95 0.00 0.00 6.64 2.61 2.84
L-R1-7B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
s1.1-32B 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
s1.1-14B 95.97 0.00 0.00 0.24 0.00 0.00 0.24 0.00 0.00 0.00 0.00 3.55
s1.1-7B 87.91 0.00 0.00 6.64 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.45
Table 9: Format types rates in greedy decoding setting.
C.1 Division of Different Format
This section presents a comprehensive taxonomy of format errors and emphasizes the importance of addressing these issues in future research endeavors. Unlike conventional LLMs, LRMs frequently exhibit format errors, manifested when responses fail to adhere to the requirement of generating a single End Thinking Marker ($\mathtt{ETM}$). This inconsistency creates significant challenges in distinguishing between the thinking process and the final answer. Format errors represent a distinct category from over-thinking and under-accuracy phenomena, yet similarly indicate vulnerability to elementary reasoning traps.
As illustrated in Table 6, we have identified and classified 12 distinct format types, each assigned a unique ID. These format types can be broadly categorized into Strict Format and Loose Format, with Strict Format representing a specialized subset of Loose Format.
Loose Format. ID-100 and ID-101 represent two standard decoding modes, categorized as Strict Format. Both contain only one $\mathtt{ETM}$. Among these, ID-100 includes a thinking process, while ID-101 only generates the $\mathtt{ETM}$ and final answer.

### v3

@@ 4.2 Evaluation Metrics
We calculate accuracy metrics under both strict and loose formatting requirements, respectively. We use GPT-4o as the evaluator to assess the correctness of the responses [^6] If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed., with the evaluation prompt in Appendix B.2. For greedy sampling, we directly calculate the accuracy rate. For top-p sampling, we utilize two metrics: Pass@1 and Acc@k. Pass@1 follows DeepSeek-R1 Guo et al. (2025), and Acc@k is the percentage of questions with all k answers correct. The two metrics use k=5, and their detailed definitions can be found in Appendix B.3. Notably, S-Corr \ L-Corr represents the upper bound for pass@1 and acc@5 under strict \ loose formatting requirements.
4.3 Main Results
We observed the following: First, state-of-the-art LRMs, such as DeepSeek-R1 and Qwen3, do not demonstrate a distinct advantage in efficiency. In contrast, Sky-T1-32B, which undergoes specific optimizations to mitigate overthinking using SimPO, achieves the highest efficiency. Second, The L-R1-DS 7B/14B/32B models are further post-trained from the DS-R1-7B/14B/32B models. The L-R1-DS models tend to produce longer responses, suggesting that while additional post-training may enhance the model’s capability for complex reasoning, it comes at the cost of response efficiency. Finally, the s1.1 models generate considerably longer responses than the DeepSeek-R1-Distilled models. Despite both models being trained solely with SFT to acquire long-COT reasoning ability, the DeepSeek-R1-Distilled models use 800K training samples, while the s1.1 models are trained on only 1K. This discrepancy suggests that the smaller training set may lead to superficial imitation of long reasoning patterns, resulting in verbose thinking on simple questions.
Several LRMs exhibit under-accuracy and limited robustness on simple questions.
First, our observations find that, despite employing deep reasoning, most LRMs tend to exhibit lower accuracy on simple questions compared to traditional LLMs. For example, DS-R1-1.5B and EXAONE-2.4B achieve just above 50% acc@k. Second, many LRMs struggle with robust correctness in top-p sampling, where acc@k is significantly lower than pass@1. This issue is particularly pronounced in smaller LRMs. For instance, DS-R1-1.5B achieved 81.47% pass@1 but only 54.50% acc@k.
5 Efficiency Analysis
5.1 Analysis across Question Types
To better understand the efficiency differences of LRMs across question types, we analyze the average response tokens across 4 main categories, 28 subcategories, and two languages. The results are displayed in Table 4 and Appendix C.2.
LRMs exhibit a substantial increase in response length across all four major categories, 28 subcategories, and two languages.
As shown in Table 4, for each of the four major categories, the average response length of LRMs exceeds that of LLMs by more than a factor of ten. Response lengths also increase significantly across all subcategories (see Appendix C.2). This suggests that while LRMs are primarily trained on reasoning data to produce long CoT style responses, this stylistic pattern generalizes well across a wide range of question types.
Moreover, 23 out of the 28 LRMs produce longer responses to questions in English than Chinese.
LRMs exhibit the most significant increase in ART for instruction following questions and tend to over-explore when the solution space is vast.
As shown in Table 4, although small LLMs provide the most concise responses to instruction following questions, LRMs generate dramatically longer outputs—96.0 times longer in English and 57.3 times longer in Chinese than small LLMs. To investigate the cause, we further analyze the subcategories of instruction following questions. As shown in Appendix C.2, average tokens is notably longer in the subcategories of length constraints, character constraints, and sentence constraints. These three question types share a similar characteristic: their correctness is easy to verify, but the solution space is vast. We find that, although the model quickly identifies a correct answer, it becomes trapped in the search space, continually exploring alternatives and failing to stop in time. A case can be seen in Table 21. This phenomenon is more pronounced in families with lower efficiency, such as s1.1 and EXAONE.
Figure 3: (a) Comparison of first round and additional token costs for each LRM. (b) Distribution of solution rounds for each LRM.
5.2 Thinking Solution Analysis
To better understand the causes of inefficiency in LRMs on S1-Bench, we analyze the solution rounds of their thinking processes [^7] We only analyze well-formatted thinking processes with correct final answers, as incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed thinking processes cannot be precisely extracted.. We first use DeepSeek-v3 to segment each thinking process into several solutions, each defined as a point at which LRMs explicitly arrives at a conclusion that matches the correct answer. We then compute the average token counts in the first solution. The detailed experimental setup is provided in Appendix C.3. Our analysis reveals the following:
The token consumed in the first solution of LRMs significantly exceeds that of validator LLMs, as shown in Figure 3 (a).
This suggests that LRMs may involve unnecessary reasoning steps in each solution, which could be one of the reasons for their inefficiency.
The primary reason for efficiency gaps between LRMs lies in the number of redundant solution rounds they generate, rather than the token cost in the initial round.
As shown in Figure 3 (a), although total thinking token counts vary widely across LRMs, their token counts in the initial round are similar and only account for a small fraction of the total. Figure 3 (b) further shows the distribution of solution rounds on S1-Bench, revealing that LRMs with longer thinking processes tend to generate more solution round, and this redundancy greatly increases computational cost. Furthermore, further experiments reveal that the redundancy in the reasoning process gradually increases over time. Appendix C.4 presents the experimental details.
6 Error Analysis
This section analyzes the errors made in the thinking process. Specifically, we utilize DeepSeek-v3 to categorize the responses of LRMs into four cases and compute the corresponding proportions: (1) Final answer correct; thinking process entirely accurate. (2) Final answer correct; thinking process contains intermediate errors. (3) Final answer incorrect; correct answer mentioned in thinking process. (4) Final answer incorrect; correct answer never mentioned in thinking process. The classification details are in Appendix C.5; results are shown in Figure 4. Key findings include:
Lower-accuracy LRMs tend to produce less reliable reasoning chains; even when they arrive at the correct final answer, their intermediate steps often contain errors (light green).
LRMs with high accuracy (e.g., DS-R1) show almost no flawed reasoning steps, whereas those with lower accuracy (e.g., DS-R1-1.5B) often generate incorrect intermediate conclusions, further indicating that they lack robust reasoning ability.
Although LRMs sometimes mention the correct answer during reasoning, they may deviate and ultimately produce incorrect final answers (light red).
In one case, the LRM initially arrived at the correct answer but undermined it through excessive verification, a case can be seen in Table 24. In another case, the LRM directly denies the correct answer, a case can be seen in Table 23.
7 Gut Moment
Figure 5: Top: Count of “gut moments” across models. Bottom: Probability of “gut moments” by question type.
[…]
First, all LRMs show the “gut moment” phenomenon to varying degrees, which is more evident in the Qwen, DeepSeek, and Light-R1 families and Hunyuan-T1. Second, LRMs show stylistic differences in expressing “gut moment.“ For example, the Qwen family often views questions as simple, whereas the DeepSeek-distilled models show more diverse difficulty comments. Third, some LRMs show significantly stronger “gut moment” in Chinese than in English, such as the Qwen and DeepSeek families, likely due to a higher proportion of Chinese in their training data. Finally, the “gut moment” is most evident in reasoning questions and rarely appears in analytical questions, except in DeepSeek-distilled models.
Figure 6: Average response tokens in the easy category vs. all samples. Dots show difference: easy minus all.
To investigate whether the early sense of a question as “easy” leads to a corresponding reduction in response length, we compare the average response tokens for questions in the easy category versus all samples. The results are shown in Figure 6. Except for L-R1-32B, other LRMs do not exhibit a noticeable decrease in response length when questions are viewed as “easy”; in fact, 21 out of 28 LRMs showed an increase in response length under this condition. This suggests a discrepancy between the LRM’s initial sense of difficulty and its generative behavior, the causes and improvements of which warrant further investigation.
8 Conclusion
This paper introduces S1-Bench, the first benchmark designed to evaluate system 1 thinking capabilities in LRMs. We conduct extensive evaluations across 28 LRMs, revealing their inefficiency, inadequate accuracy, and limited robustness when handling simple questions. Additionally, we observe “gut moment” and find a gap between their difficulty perception and generation length. Overall, this work paves the way toward dual-system compatibility in the development of LRMs.
Limitations
Although S1-Bench pioneers the evaluation of system 1 thinking in LRMs, it still has several limitations. First, due to our emphasis on ensuring the uniqueness of each sample during dataset construction—for instance, including only one question for basic arithmetic operations such as addition, subtraction, and multiplication—the overall scale of the benchmark remains limited. As a next step, we plan to expand the scale of S1-Bench. Second, while recent months have seen a surge in newly released open-source LRMs, we have only evaluated 28 representative models and have not covered the full spectrum of available models. Lastly, we do not propose methods to improve the efficiency of LRMs on system 1 tasks in this work; this will be the focus of our future research.
Appendix A More Information of S1-Bench Construction
A.1 Benchmark Statistics
We survey studies on improving the efficiency of LRMs, as there is potential overlap between these studies and the technical approaches aimed at enhancing system 1 thinking in LRMs. Table 7 presents the results of our survey. We compile the benchmarks used in these studies for evaluation, that are typically used to verify whether models achieve efficiency improvements. Benchmarks that appear more than four times include: MATH500 Hendrycks et al. (2021b), GSM8K Cobbe et al. (2021), AIME24/25 MAA Committees , GPQA Rein et al. (2024), AMC23 AI-MO (2024), MMLU Hendrycks et al. (2021a), Olympiad-Bench He et al. (2024), SVAMP Patel et al. (2021), LiveCodeBench Jain et al. (2025), and CommonSenseQA Talmor et al. (2019).
The accuracy shown in Table 1 is the average result of the four models, Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B, at temperature 0, using GPT-4o as the evaluator.
A.2 Subcategories in S1-Bench
Figure 7 shows the pie chart distribution of 28 subcategories in S1-Bench. For more details on the subcategories, please refer to Table 8,9.
A.3 Prompt for S1-Bench construction
This section presents the prompts used in the construction of S1-Bench, including the Initial Generation prompt, the Discriminating Generation Quality prompt, and the Reduce Difficulty prompt. See Table 10 for details.
Appendix B Baseline Models and Evaluation Details
B.1 Baseline Model Details
Table 11 presents the abbreviations, IDs, and URLs of LLMs used in this paper. Table 12 displays the abbreviations, IDs, URLs, organizations, training algorithms, and training data volumes of open-source LRMs evaluated in this study.
B.2 GPT-4o and Human Evaluation
We use GPT-4o as the evaluator to assess the correctness of the responses. If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed. The evaluation prompt is provided in Table 13.
To evaluate the consistency between the GPT-4 judge’s assessments and human judgments, we conduct a comprehensive human evaluation study involving three of the authors. Specifically, we randomly sample 20 question-answer pairs from each model’s greedy decoding results, resulting in a dataset of 640 pairs derived from 32 models (including 4 verifier LLMs and 28 LRMs). The questions, reference answers, and model responses are then presented to three annotators, who independently judge the correctness of each model response. The final human evaluation results are determined through majority voting. Ultimately, the Cohen’s Kappa between the human evaluators and the GPT-4 judge is calculated to be 0.83, indicating an exceptionally high level of agreement.
B.3 Accuracy Metrics Details
@@
B.4 Types and Analysis of Format Errors
This section introduces a comprehensive taxonomy of format errors and highlights the importance of addressing these issues in future research. Unlike conventional LLMs, LRMs frequently exhibit format errors. These errors are defined by failing to use a unique end thinking marker (e.g., </think>) to separate the thinking process from the final answer. Format errors increase the difficulty of distinguishing the thinking process from the final answer and reveal the vulnerability of LRMs in following predefined formats.
To illustrate this phenomenon, we identify 12 distinct types of response formats produced by LRMs, each assigned a unique ID, as shown in Table 5. These 12 types are further grouped into three major categories:
  • Standard-Conforming Responses: These responses meet the expected format by including exactly one end thinking marker (e.g., </think>) to delimit the thinking process from the final answer. Among these, type ID-100 includes a thinking process, while ID-101 omits it. The proportion of such responses is measured using the S-Corr metric.
  • Unreadable Responses: These refer to generation failures, including cases where LRMs produce endlessly thinking content or solely produce end thinking markers. The proportion of all other (i.e., readable) responses is measured using the L-Corr metric.
[…]
Table 14 and Table 15 present the distributions of 12 format types under top-p sampling and greedy sampling, respectively. we find: (1) The infinite generation phenomenon is widespread across most LRMs, particularly concentrated in LRMs with fewer than 32B parameters. (2) The Nemotron and EXAONE families frequently produce correctly formatted responses without any explicit thinking processes. This behavior can be viewed as a mechanism for mitigating over-thinking. However, the EXAONE family still exhibits substantial overthinking tendencies, suggesting that LRMs’ capability to respond without visible reasoning and their tendency to overthink may be orthogonal characteristics. (3) None of the evaluated LRMs exhibited behaviors classified as ID-205/206.
@@ B.4 Types and Analysis of Format Errors
(standard) (unmatched) (number) process answer
Standard 100 $\surd$ 1 $\surd$ $\surd$
@@
Readable but Malformed 200 $\surd$ 1 $\surd$ $\times$
@@
Unreadable 300 $\surd$ $\surd$ $\geq$1 $\times$ $\times$
@@
Table 5: Twelve types of response format.
Appendix C More Experimental Setups & Results
C.1 Greedy Sampling Results
Table 16 presents the performance of LRMs on S1-Bench under greedy sampling. While overall accuracy improves compared to top-p sampling, issues of inefficiency and accuracy degradation on simple questions remain.
C.2 Efficiency Analysis across Subcategories.
Figure 9 illustrates the average response tokens across the 28 subcategories. In the heatmap, both models (rows) and subcategories (columns) are ordered in descending order according to their average number of response tokens.
C.3 Solution Analysis Details
For solution analysis, We only use well-formatted thinking processes with correct final answers, as incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed thinking processes cannot be precisely extracted. The segmentation process is performed by DeepSeek-v3, with prompts detailed in Table 17. We compute the average token count in the first solution round; if no solution is found, we use the token count of the entire thinking process.
C.4 Thinking Redundancy Analysis
We conduct a similarity analysis to analyze how information redundancy in the thinking processes changes as reasoning sequences increase. Specifically, we first divide the complete thinking process into k equal-length segments [^10] We set k=15, changing its value does not affect the conclusions.. Then, we encode each segment using the all-MiniLM-L6-v2 model [^11] https://huggingface.co/sentence-transformers/allMiniLM-L6-v2. For each segment, we calculate the cosine similarity with all its preceding segments and use the maximum similarity as a measure of its information redundancy. As shown in Figure 10, information redundancy increases across all four main categories as reasoning sequences increase. Sky-T1-32B shows overall lower similarity, which stems from its shorter thinking process, but still demonstrates an upward trend.
C.5 Error Analysis Details
In error analysis, we only use well-formatted samples, as malformed thinking processes cannot be precisely extracted. For samples with correct final answers, we categorize them based on whether the thinking process contains explicit incorrect conclusions in intermediate steps. For samples with incorrect final answers, we categorize them based on whether the correct answer is mentioned at least once during reasoning. We use DeepSeek-v3 for categorization, with prompts provided in Table 18.
C.6 Gut Moment Analysis Details